OpenStudy (anonymous):

3. Solve the system by graphing: y = 1/2x + 9 y = 1/2x - 1 4.What is the solution to the following system? y = 4x + 3 y - 6 = x 5. Is the coordinate (1, 2) a solution to the system? y = x + 1 2x - y = 0 6. True/False: When you are looking at a graph of a system, the solution is the coordinate where the two lines cross.

OpenStudy (texaschic101):

4. y = 4x + 3 now sub 4x + 3 in for y in the other equation y - 6 = x 4x + 3 - 6 = x 4x - x = 6 - 3 3x = 3 x = 1 now sub 1 in for x in either equation to find y y - 6 = x y - 6 = 1 y = 6 + 1 y = 7 ================ 5. (1,2) y = x + 1 2 = 1 + 1 2 = 2 (correct) 2x - y = 0 2(1) - 2 = 0 2 - 2 = 0 0 = 0 (correct) yes...(1,2) is not a solution ================== 6. true...the solution is where the two lines intersect as for question 3.....do you know how to graph ?

OpenStudy (texaschic101):

oops...for # 5....it is a solution

OpenStudy (texaschic101):

3. y = 1/2x + 9 y intercept is (0,9) x intercept can be found by subbing in 0 for y 0 = 1/2x + 9 -1/2x = 9 x = 9/(-1/2) x = 9 * - 2 x = - 18 plot your points (0,9) and (-18,0) start at (-18,0)...and since the slope is 1/2....go up 1, and to the right 2, then up one, and to the right 2 and you should cross the y axis at (0,9) y = 1/2x - 1 y intercept is (0,-1) x intercept = 0 = 1/2x - 1 -1/2x = -1 x = -1/(-1/2) x = -1 * -2 x = 2 x intercept is (2,0) plot your points (0,-1) and (2,0) start at (0,-1) ...and since slope is 1/2...go up 1 and to the right 2, then up 1, and to the right 2, and you should cross the x axis at (2,0) Once you have graphed these lines, you will realize that they are parallel lines, and that there is no solution because the lines do not intersect